TheSmartOne (thesmartone):

An electron in a hydrogen atom in the n = 6 energy level emits 109.4 kJ/mol of energy in a transition to a lower energy level. To what energy level does the electron fall?

TheSmartOne (thesmartone):

First question, nice

OpenStudy (melissa_something):

hi, do you still need help?

TheSmartOne (thesmartone):

Yes

TheSmartOne (thesmartone):

I've tried Rydberg's equation and also Bohr's but no idea \(\Large \frac{1}{\lambda}=R\left(\frac{1}{n^{2}_{f}} - \frac{1}{n^{2}_{i}}\right)\) After using \(\Large \Delta E = \frac{hc}{\lambda}\) And the other formula \(\Large \Delta E = -b\left(\frac{1}{n^{2}_{f}} - \frac{1}{n^{2}_{i}}\right)\) where b = 2.178 * 10^-18 J h = 6.63 * 10^-34 J s c = 2.998 * 10^8 m s^-1 R = 1.0974 x 107 m^-1

TheSmartOne (thesmartone):

\(\Large \Delta E = -b\left(\frac{1}{n^{2}_{f}} - \frac{1}{n^{2}_{i}}\right)\) \( 109.4 ~kJ~mol^{-1}= -2.178\cdot 10^{-18} J\left(\frac{1}{n^{2}_{f}} - \frac{1}{6^2}\right)\) \(\Large \frac{109400 ~J~mol^{-1}}{ -2.178\cdot 10^{-18} J}=\left(\frac{1}{n^{2}_{f}} - \frac{1}{6^2}\right)\) \(\Large -5.02295684e22~ mol^{-1} + \frac{1}{36}=\frac{1}{n^{2}_{f}} \) like idek what I'm doing x'd

OpenStudy (lowkey.s):

The formula for the energy is E(n) = -13.6 eV * (1/n^2) E(n) - E(6) = 13.6 eV * (1/n^2 - 1/6^2) = 109.4 kJ /mol 109.4 kJ / mol = 109.4 * (10^3 J / kJ) * (1 eV / 1.6 * 10^-19 J) * (1 mol / 6.02 *10^23 e) * (1 e / atom) = 1.14 eV n = [1.14 eV / 13.6 eV + 1/6^2]^(-1/2) = 2.99... ~ 3. The answer is n = 3.

OpenStudy (lowkey.s):

Dang that was long! LOL cxxx

TheSmartOne (thesmartone):

LowKey, I don't need a copy pasted answer from some other source. Can you explain where you got the formula E(n) = -13.6 eV * (1/n^2) from?

OpenStudy (jiteshmeghwal9):

Well niels Bohr derived the formula for hydrogen like species \[E(n)=-13.6 \frac{z^2}{n^2}eV\]